OpenStudy (anonymous):

How can you write the expression with a rationalized denominator? √3-√6/√3+√6

OpenStudy (anonymous):

to rationalise the denominator u need to multiply both num and denom by the conjugate of the denom

OpenStudy (anonymous):

multiply bottom and up in sqrt3-sqrt6

OpenStudy (anonymous):

for √3+√6 the conjugate is √3- √6 hence u have to multiply with √3+√6 (both num and denom) (√3-√6)*(√3-√6)/(√3+√6)*(√3-√6)= (√3-√6)^2/(3-6) =(3+6 -2√18)/(-3)= = (9-2*3√2)/(-3) =-3+2√2

OpenStudy (anonymous):

Thank you sooo much!!!! :)

OpenStudy (anonymous):

welcome